Analyze the data and infer the rate order of each reaction with respect to each reactant. Clearly express the rate law for Reaction 1 and Reaction 2. Calculate the rate constant (with the correct units) for Reaction 1 and 2. Please show all work in a highly organized, very legible, easy to follow format.
Answer #1

Reaction 1 CH3CH2 BO + H2O CH3 CH2OH + HBO So kao rate = k [CH₃ CH₂ Br] [#₂0] karate constant for 8 - Trial 1, -2.11x10 ....Reaction (2) (CH3) 3 C Br + H₂O - (CH3) 3 COH + HBD Trrial 1 mi = kx (1,51x172] [1.56x1523° = 3.89x1o5...) Troial 2 = 6.78*10